light can have a wavelength of 550mm the energy of a photon of light is??
Answer #1

Given that wavelength of light is 550 mm light wavelength is in nm hence 550 mm = 5.5 *10^8 nm = 0.55 m

Now, formula for energy

E= hc/ lambda

Where h is planks constant = 6.626 *10^ -34

C is speed of light 3*10^8 m/s

On Putting values we get,

E = 6.626*10^-34*3*10^8/0.55

E = 36.142*10^-26 J

If lambda is in nm then answer is as follws 550*10^-9

E= 6.626*10^-34*3*10^8/550*10^-9

E= 6.57*10^-19 J

I give two answer one is using 550 mm and one is using 550 nm as shown above